OpenStudy (anonymous):

5(x + 2) greater than or equal to 7x + 2(x − 1) I need help

OpenStudy (alekos):

so I guess you need to solve for x, is that right?

OpenStudy (anonymous):

yes can you please help me?

OpenStudy (alekos):

\[5(x+2) \ge 7x+2(x-1)\]

OpenStudy (alekos):

so lets start with the LHS and expand

OpenStudy (anonymous):

ok so we distribute first right

OpenStudy (alekos):

yep

OpenStudy (alekos):

over to you

OpenStudy (anonymous):

so it will be 5x and 2x

OpenStudy (anonymous):

5(x + 2) greater than or equal to 7x + 2(x − 1) 5x + 10

OpenStudy (anonymous):

7x + 2x-2

OpenStudy (anonymous):

5x +10 < 9x-2

OpenStudy (alekos):

NO \[5x+10\ge9x-2\]

OpenStudy (alekos):

Now you'll need to get x's on the left and constants on the right

OpenStudy (anonymous):

@smart103 What was the correct answer?

OpenStudy (alekos):

Next step is \[5x-9x \ge -2-10\]

OpenStudy (alekos):

then we have \[-4x \ge-12\]

OpenStudy (alekos):

multiply both sides by -1 \[4x \le12\]

OpenStudy (alekos):

finally divide both sides by 4 \[x \le3\]